### v2

# Hybrid Bernstein Normalizing Flows for Flexible Multivariate Density Regression with Interpretable Marginals

Abstract
Density regression models allow a comprehensive understanding of data by modeling the complete conditional probability distribution. While flexible estimation approaches such as normalizing flows (NFs) work particularly well in multiple dimensions, interpreting the input-output relationship of such models is often difficult, due to the black-box character of deep learning models. In contrast, existing statistical methods for multivariate outcomes such as multivariate conditional transformation models (MCTMs) are restricted in flexibility and are often not expressive enough to represent complex multivariate probability distributions. In this paper, we combine MCTMs with state-of-the-art and autoregressive NFs to leverage the transparency of MCTMs for modeling interpretable feature effects on the marginal distributions in the first step and the flexibility of neural-network-based NF techniques to account for complex and non-linear relationships in the joint data distribution. We demonstrate our method’s versatility in various numerical experiments and compare it with MCTMs and other NF models on both simulated and real-world data.
1 Introduction
Many real-world regression problems involve modeling of complex conditional distributions of an outcome $Y$ given an input vector $\mathbf{x}=[x_{1},\ldots,x_{D}]^{\top}$. In such instances, a mere prediction of the mean would be insufficient to capture the aleatoric uncertainty inherent in the data-generating process and a comprehensive understanding requires estimating the conditional distribution $Y|\mathbf{x}$ (Kneib2021). This becomes more challenging for multivariate outcomes $\mathbf{Y}\in\mathbb{R}^{D}$, both in terms of estimation and model interpretation. While normalizing flow (NF(Papamakarios2018) models can represent such complex conditional multivariate distributions, their deep neural network architectures often hinder interpretability. Statistical methods for flexible density estimation in high dimensions have largely focused on graphical models, often involving copulas (e.g., Liu2011; Bauer2016; Nagler2016). A notable exception is the MCTM, a transformation model approach (Hothorn2018) that shares similarities with NFs and models the joint relationship similar to copulas. While less flexible than NFs, MCTMs preserve the interpretability of the feature-outcome relationship.
Figure 1: An MCTM (orange) applied to a complex data distribution (blue). The model captures marginals well but fails to capture the dependence structure.
Our Contribution
To bridge the gap between flexible black-box and rigid but interpretable statistical models, we propose a hybrid approach combining the transparency of MCTMs with the flexibility of NFs. We thereby provide a method that can be applied if the understanding of feature effects on each response variable is indispensable. At the same time, our method allows complex modeling of the dependency structure of the outcome dimensions using the idea of masking from autoregressive flows. We evaluate its effectiveness against MCTMs and other density estimation models on simulated and real-world datasets.
Notation
We refer to random variables with a capital Latin letter, e.g., $Y$, and their realizations (observed or sampled) with a lowercase Latin letter, e.g. $y$. The corresponding CDF is referred to as $F_{Y}$ and the probability density function (PDF) as $f_{Y}$. For sets and vectors, we use bold-faced variables, e.g. $\mathbf{Y}=\{Y_{1},\dots,Y_{J}\}$. For individual parameters, optimized during the training procedure we use lowercase Greek letters, e.g. $\vartheta$, and bold-faced Greek letters for sets or vectors of parameters, e.g. $\bm{\vartheta}=\left(\vartheta_{1},\ldots,\vartheta_{M}\right)^{\top}$. Bold capital Greek letters indicate tuples, e.g. $\bm{\Theta}=\left(\bm{\theta}_{1},\bm{\theta}_{2},\alpha,\beta\right)$.
2 Background and Related Work
Several methods have emerged for modeling complex conditional distributions. Among these methods, transformation-based models have proven to be highly flexible and have been extensively used since their introduction by Box1964.
2.1 Probability Transformation
Transformation models are based on the probability-transformation theorem, using a bijective and continuously differentiable transformation function $h$ to map a complex distribution $F_{Y}$ to a simpler one $F_{Z}$ (often standard normal) (Kneib2021) via $F_{Y}=F_{Z}(h(y)).$ This allows density estimation without restrictive shape assumptions (Hothorn2018). The density $f_{Y}$ is calculated from the base density $f_{Z}$ by
$f_{Y}(y)=f_{Z}\left(h(y)\right)\left|\det\nabla{h}(y)\right|,$ (1)
where the absolute Jacobian determinant $\left|\det\nabla{h}(y)\right|$ reflects the change in density induced by the transformation. To sample from $F_{Y}$, we can draw samples $z$ from $F_{Z}$ and apply the inverse transformation $h^{-1}$ to obtain $y$.
2.2 Normalizing Flows and Transformation Models
NFs apply the probability-transformation theorem using a series of $K$ simple bijective transformation functions $h_{k}$ to form more expressive transformations $h(z)=h_{K}\circ h_{K-1}\circ\dots\circ h_{1}(z)$ (see, e.g. Papamakarios2021, for a comprehensive review). Other methods use a single flexible transformation like sum-of-squares polynomials (Jaini2019) or splines (Durkan2019). Conditional transformation models (CTMs) also focus on single transformation function, which, however, can be easily interpreted (see, e.g., Hothorn2018). This idea has also recently been combined with neural-networks (Baumann2021; Sick2021; Arpogaus2023; Kook2024).
A fundamental difference between NFs and CTMs is the definition of the transformation direction (Kook2024): NFs transform a simple base density $f_{Z}(z)$ into a more complex target density $f_{Y}(y)$ through a series of transformations. CTMs do it vice-versa, using only a single transformation function, often utilizing flexible Bernstein polynomials (see Section A.1 for details).
CTMs allow interpretability through the use of structured additive predictors (SAPs) (Klein2022; Rugamer2023) (see Sections A.5 and B.3). The black-box character of neural networks in NFs hinders interpretability, and variants using decision trees (Papastefanopoulos2025) and graphical models (Wehenkel2021a) have been explored in recent years.
The previous concepts can be further extended to multivariate objects $\mathbf{y}=(y_{1},\ldots,y_{J})^{\top}$ and can also be applied conditional on $U$ features $\mathbf{x}=(x_{1},\ldots,x_{U})^{\top}$ as described in the following.
2.3 Multivariate and Conditional Models
To model a potentially multivariate conditional distribution $F_{\mathbf{Y}|\mathbf{X}=\mathbf{x}}$, we define $z_{j}=h(y_{j},\bm{\theta}_{j})\text{ with }\bm{\theta}_{j}=c_{j}(\mathbf{y},\mathbf{x})$ where $h$ is a bijective transformation function for the element $y_{j}$ with parameters $\bm{\theta}_{j}$ obtained from the conditioner $c_{j}$ (Papamakarios2021). The latter can be an arbitrarily complex function that depends on the features $\mathbf{x}$ and, in the multivariate case, on some variables of the response $\mathbf{y}$. The limiting factor here is the computational tractability of the Jacobian in Equation 1.
Multivariate Conditional Transformation Models
To make computations tractable and provide a model that is better to interpret, MCTMs define a triangular transformation $H(\bm{y},\bm{\Theta}(\bm{x}))=\left(h_{1}(y_{1}|\bm{\theta}_{1}),\ldots,h_{J}(y_{J}|y_{1},\ldots,y_{J-1},\bm{\theta}_{1},\dots,\bm{\theta}_{J})\right)^{\top}$. This Tranformation is characterized by a set of linearly combined marginal basis transformations $\tilde{h}_{j}(y_{j}|\bm{\theta}_{j})$, scaled with a $(J\times J)$ triangular coefficient matrix $\bm{\Lambda}$ to encode structural information: $h_{j}(y_{j}|y_{1},\ldots,y_{j-1},\bm{\theta}_{1},\dots,\bm{\theta}_{j})=\lambda_{j,1}\tilde{h}_{1}(y_{1}|\bm{\theta}_{1})+\ldots+\lambda_{j,j-1}\tilde{h}_{j-1}(y_{j-1}|\bm{\theta}_{j-1})+\tilde{h}_{j}(y_{j}|\bm{\theta}_{j})$. The function $\bm{\Theta}:\bm{x}\mapsto(\bm{\Lambda},\bm{\theta}_{1},\ldots,\bm{\theta}_{J})$ accounts for feature effects on the parameters (Klein2022). Hence, MCTMs can only model linear dependencies between the response variables $y_{j}$, making them an inadequate choice for complex joint distributions as shown in Figure 1. This limitation motivates the development of more flexible approaches, such as the hybrid model proposed in this paper.
Normalizing Flows
In the NF literature, various approaches are used to deal with computations involving the aforementioned Jacobian. Among these, autoregressive models are one of the most prevalent ones. Due to the autoregressive structure, these NFs yield a triangular Jacobian and thus the determinant simplifies to the product of the Jacobian’s diagonal entries (Papamakarios2021; Kobyzev2021). These models factorize a multivariate distribution based on the chain rule of probability. They are typically implemented by conditioning the transformation $h_{j}$ of the $j$th element of $\mathbf{y}$ on all previous elements $\mathbf{y}_{<j}=(y_{1},\dots,y_{j-1})$ or a subset of those. This can be seen as a non-linear generalization of the triangular coefficient matrix $\Lambda$ used in MCTMs (Kobyzev2021). There are two widely adopted neural-network architectures for implementing an autoregressive conditioner $c_{j}$s (Papamakarios2018) and CFs (Dinh2017).
MAFs generalize this idea by implementing the conditioner with neural networks that incorporate autoregressive constraints through parameter masking, inspired by the masked autoencoder for distribution estimation (MADE) architecture (Germain2015). The main advantage is that the parameters can be obtained in one neural network pass $\left(\bm{\theta}_{1},\ldots,\bm{\theta}_{J}\right)^{\top}=\left(c_{1}(),c_{2}(y_{1}),\ldots,c_{J}(\mathbf{y}_{<j})\right)^{\top}$. Further, if $h$ and the conditioning network are expressive enough, they remain universal approximators to transform between any two distributions (Papamakarios2021). In practice, however, the results highly depend on the ordering of the input variables. Because some orderings can be extremely difficult to learn, multiple MAFs layers are often stacked with permutations between.
CFs can provide fast sampling and density evaluation by splitting the response vector $\mathbf{y}$ into two subsets
$\mathbf{y}=(\underbrace{y_{1},\ldots,y_{j}}_{\mathbf{y}_{\leq j}},\underbrace{y_{j+1},\ldots,y_{J}}_{\mathbf{y}_{>j}})^{\top}$
and then applying the transformation $h$ only to one subset conditioned on the other. Typically, these subsets are chosen to contain half of the variables and the result of the transformation on the first subset is then permuted and another transformation is applied to the remaining subset: $\mathbf{z}=\left(h(\mathbf{y}_{\leq d},c_{1}(\mathbf{y}_{>d},\mathbf{x}),h(\mathbf{y}_{>d},c_{2}(\mathbf{y}_{\leq d},\mathbf{x})\right)^{\top}$.
Graphical Transformation Models
are an extension of MCTMs, closely related to the methodology proposed in this paper. Herp2025 replaced the Gaussian copula by parameterizing the $\lambda_{i,j}$ entries of $\bm{\Lambda}$ using B-splines, dependent on the previous variables, resulting in an additive CF: $h_{j}(y_{j}|y_{1},\ldots,y_{j-1},\bm{\theta}_{1},\dots,\bm{\theta}_{j})=\lambda_{j,1}\left(\tilde{h}_{1}(y_{1}|\bm{\theta}_{1})\right)+\ldots+\lambda_{j,j-1}\left(\tilde{h}_{j-1}(y_{j-1}|\bm{\theta}_{j-1})\right)+\tilde{h}_{j}(y_{j}|\bm{\theta}_{j})$. To enhance expressiveness of the dependencies while preserving marginals, multiple CFs are chained via permutations. P-spline penalties mediate between the flexible CFs and the baseline MCTM. This framework improves interpretability of nonlinear conditional dependencies through local conditional pseudo-correlations and enables sparse undirected graphical models via a LASSO penalty promoting conditional independence. However, it offers less flexibility as a NF for decorrelation than the approach proposed here and has not yet been extended to the conditional setting.
Further background is given in Section A.3 and A.4.
3 Hybrid Conditional Masked Autoregressive Bernstein Flows
To bridge the gap between the limited expressiveness of MCTMs and the black-box nature of general NFs, we propose a combined approach that leverages the strengths of both methodologies. This hybrid approach allows us to capture complex dependencies in the joint distribution while retaining interpretability for the marginal distributions.
3.1 Model Specification
Let $\mathbf{y}=(y_{1},\dots,y_{J})^{\top}\in\mathbb{R}^{J}$ be a $J$-dimensional response vector and $\mathbf{x}=(x_{1},\dots,x_{U})^{\top}\in\mathbb{R}^{U}$ be a vector of $U$ features. Our goal is to learn the conditional joint density $f_{\mathbf{Y}|\mathbf{X}}(\mathbf{y}|\mathbf{x})$.
Step 1: Modeling Marginal Distributions
First, we model the marginal distributions of each response variable $Y_{j}$ given $\mathbf{X}$ using a transformation model:
$H_{1}(\mathbf{y},\mathbf{\Theta}\left(\mathbf{x}\right))$ $=\left(h_{1}(y_{1},\bm{\theta}_{1,\mathbf{x}}),\ldots,h_{1}(y_{J},\bm{\theta}_{J,\mathbf{x}})\right)^{\top}=$
$=\mathbf{w}=(w_{1},\ldots,w_{J})^{\top},$
where each element of $\mathbf{W}=(W_{1},\dots,W_{J})^{\top}$ follows the base distribution $F_{Z}$ (Figure 2).
Figure 2: The first transformation $H_{1}$ maps the marginals to the base distribution $F_{Z}$.
$H_{1}(\mathbf{y}|\mathbf{x})$
$H^{-1}_{1}(\mathbf{w}|\mathbf{x})$
Specifically, for each $j=1,\dots,J$, we assume:
$\mathbb{P}(Y_{j}\leq y_{j}|\mathbf{X}=\mathbf{x})=F_{Y_{j}|\mathbf{X}}(y_{j}|\mathbf{x})=F_{Z}(h_{1}(y_{j},\bm{\theta}_{j,\mathbf{x}})),$
where $F_{Z}$ is a known base distribution and $h_{1}(\cdot)\colon\mathbb{R}\rightarrow\mathbb{R}$ is a strictly monotonic marginal transformation function, with potentially feature-dependent parameters $\bm{\theta}_{j,\mathbf{x}}$ obtained form the conditioning function $\mathbf{\Theta}\left(\mathbf{x}\right)$.
For interpretability, a (shifted) Bernstein polynomial can be used:
$\begin{split}h_{1}(y_{j},\bm{\theta}_{j,\mathbf{x}})&=\bm{\alpha}_{j}(y_{j})^{\top}\bm{\vartheta}_{j}+\beta_{j,\mathbf{x}},\\ \bm{\alpha}_{j}(y_{j})^{\top}\bm{\vartheta}_{j}&=\frac{1}{M+1}\sum_{i=0}^{M}\text{Be}_{i}(\tilde{y}_{j})\vartheta_{ji},\end{split}$
where $\text{Be}_{m}(\tilde{y}_{j})$ is the density of a Beta distribution with parameters $i+1$ and $M-i+1$ evaluated at the normalized response $\tilde{y}_{j}=(y_{j}-l_{j})/(u_{j}-l_{j})\in[0,1]$, with $u_{j}>l_{j}$ defining the support of $Y_{j}$ [^1] The transformation is generally unbounded, as we apply linear extrapolation outside the bound of the polynomial, as all our experiments show. However, other transformations may be considered if expressiveness in the outer tails is a requirement. The vector $\bm{\vartheta}_{j}=(\vartheta_{j0},\dots,\vartheta_{jM})^{\top}$ contains the Bernstein coefficients, which are constrained to be increasing for monotonicity (see Section A.1 for details). The shift term $\beta_{j,\mathbf{x}}$ allows the marginal distribution of $Y_{j}$ to vary with the features. Instead of a feature-dependent shift term, we can also let $\bm{\vartheta}_{j}$ depend on $\mathbf{x}$ to change the shape of the transformation $h_{1}$ (or combine feature-dependent shift with feature-dependent $\bm{\vartheta}_{j}$s).
Step 2: Modeling Dependency Structures
We model the dependencies between elements of $\mathbf{W}$ using an autoregressive flow $H_{2}(\mathbf{w},\mathbf{\Psi}(\mathbf{w},\mathbf{x})):\mathbf{W}\to\mathbf{Z},\mathbf{w}\mapsto\mathbf{z}$:
$\begin{split}z_{1}&=h_{2}(w_{1}|\bm{\psi}_{1,\mathbf{x}}),\\ z_{j}&=h_{2}(w_{j}|\bm{\psi}_{j,\mathbf{w}_{<j},\mathbf{x}}),\quad j=2,\dots,J,\end{split}$
where $h_{2}(\cdot)$ is an increasing transformation function, with parameters $\bm{\psi}_{j,\mathbf{w}_{<j},\mathbf{x}}$ depending on previous elements of $\mathbf{w}$ and features $\mathbf{x}$.
Figure 3: The second transformation $H_{2}$ removes the dependency structure.
$H_{2}(\mathbf{w}|\mathbf{x})$
$H^{-1}_{2}(\mathbf{z}|\mathbf{x})$
For $h_{2}$, flexible transformation functions like Bernstein polynomials or splines can be used.
We combine ideas from CFs and MAFs to construct the conditioner. Similar to CFs, we do not apply any transformation to $w_{1}$ but pass it as a conditional input to the masked neural network that estimates the parameters of the subsequent transformations $\bm{\Psi}(\mathbf{w},\mathbf{x})=\left(\bm{\psi}_{\mathbf{x}},\bm{\psi}_{\mathbf{w}_{<1},\mathbf{x}},\ldots,\bm{\psi}_{\mathbf{w}_{<J},\mathbf{x}}\right)^{\top}$. Binary masking, similar to MADE, ensures that the $j$th parameters do not depend on $\mathbf{z}_{1,\geq j}$. Autoregressive transformations can be stacked if a single transformation is not expressive enough. Through this autoregressive transformation, we aim to “de-correlate” the elements of $\mathbf{W}$ such that $\mathbf{Z}$ follows a multivariate standard normal distribution $\mathcal{N}_{J}(\mathbf{0},\mathbf{I})$.
The joint density of the original response vector $\mathbf{Y}$ is then: [^2] We use the short notation $H_{1}\left(\mathbf{y},\mathbf{\Theta}\left(\mathbf{x}\right)\right)=H_{1}(\mathbf{y}|\mathbf{x})$ and $H_{2}\left(\mathbf{w},\mathbf{\Psi}\left(\mathbf{w},\mathbf{x}\right)\right)=H_{2}(\mathbf{w}|\mathbf{x})$.
$f_{\mathbf{Y}|\mathbf{X}}\left(\mathbf{y}|\mathbf{x}\right)$ $=f_{Z}\left(H\left(\mathbf{y}|\mathbf{x}\right)\right)\left|\nabla H\left(\mathbf{y}|\mathbf{x}\right)\right|$
$=f_{Z}\left(H_{2}\left(H_{1}\left(\mathbf{y}|\mathbf{x}\right)|\mathbf{x}\right)\right)$
$\quad\cdot\left|\nabla H_{2}\left(H_{1}\left(\mathbf{y}|\mathbf{x}\right)|\mathbf{x}\right)\nabla H_{1}\left(\mathbf{y}|\mathbf{x}\right)\right|.$
Model Training and Inference
We train the hybrid model by minimizing the conditional negative log-likelihood of the data $\mathcal{D}$,
$\operatorname*{NLL}(\bm{\omega})$ $=-\sum_{(\mathbf{y},\mathbf{x})\in\mathcal{D}}\log f_{\mathbf{Y}|\mathbf{X}}\left(\mathbf{y}|\mathbf{x}\right)$
$=-\sum_{(\mathbf{y},\mathbf{x})\in\mathcal{D}}\log f_{Z}\left(H\left(\mathbf{y}|\mathbf{x},\bm{\omega}\right)\right)\left|\nabla H\left(\mathbf{y}|\mathbf{x},\bm{\omega}\right)\right|,$
to optimize the weights $\bm{\omega}$ of the conditioning functions $\bm{\Theta}_{x}$ and $\bm{\Psi}_{x}$ parameterizing the normalizing flow $H\left(\mathbf{y}|\mathbf{x},\bm{\omega}\right)=H_{2}\circ H_{1}(\mathbf{y}|\mathbf{x},\bm{\omega})$.
This involves optimizing the parameters of the marginal transformation functions $H_{1}$ and the parameters of the autoregressive flow $H_{2}$, which are controlled by masked neural networks. Optimization can be done using any suitable gradient-based optimization algorithm, such as Adam (Kingma2017a). Once the model is trained, we can sample from the joint distribution of $\mathbf{Y}|\mathbf{X}$ by:
  1. 1.
    Sample $\mathbf{z}$ from the base distribution $F_{Z}$.
  2. 2.
    Apply the inverse autoregressive flow to obtain $\mathbf{w}=H_{2}^{-1}(\mathbf{z}|\mathbf{x})$.
  3. 3.
    Apply the inverse marginal transformation to obtain $\mathbf{y}=H_{1}^{-1}(\mathbf{w}|\mathbf{x})$.
In short: $\mathbf{y}=H_{1}^{-1}(H_{2}^{-1}(\mathbf{z}|\mathbf{x})|\mathbf{x})$ with $\mathbf{z}\sim F_{Z}$.
3.2 Interpretability
Our hybrid approach combines CTM’s interpretability with the flexibility of autoregressive NFs: The marginal transformation step ($H_{1}$) may use shifted Bernstein polynomials, which allows quantifying feature effects on the marginal distribution, similar to coefficients in linear models or generalized additive models (GAMs) (Wood2017). For instance, with a logistic base distribution, linear effect coefficients in the SAP can be interpreted as linear changes on the level of log-odds ratios (see Sections A.5 and B.3 for more details). Our model prioritizes marginal interpretability while accepting a trade-off for the dependence structure to gain flexibility via the masked autoregressive flow in step two ($H_{2}$). It is suitable for scenarios requiring understanding individual feature effects while modeling complex relationships between response variables. Understanding how features affect the dependence structure in this step is challenging and presents another open research question.
3.3 Relation to Copula Methods
(a) Original Data
(b) Normalized Marginals
(c) Uniform Marginals
Figure 4: Illustration of the hybrid approach on the Moons dataset. (a) The original data exhibits a non-linear dependency structure. (b) After applying $H_{1}$, the marginal distributions follow a normal distribution. (c) The autoregressive flow $H_{2}$ further transforms the data to obtain approximately independent uniform marginals, implicitly modeling the copula function.
Copulas model multivariate distributions by separating marginal distributions from the dependence structure. Sklar’s Theorem states that a copula function can express any CDF. Let $F_{\mathbf{Y}|\mathbf{X}}(\mathbf{y}|\mathbf{x})$ be the joint CDF of the response vector $\mathbf{Y}=(Y_{1},\dots,Y_{J})^{\top}$ given features $\mathbf{X}$. Sklar’s theorem implies a copula function $C(u_{1},\dots,u_{J}|\mathbf{x})$ such that:
$F_{\mathbf{Y}|\mathbf{X}}(y_{1},\dots,y_{J}|\mathbf{x})$
$=C(F_{Y_{1}|\mathbf{X}}(y_{1}|\mathbf{x}),\dots,F_{Y_{J}|\mathbf{X}}(y_{J}|\mathbf{x})|\mathbf{x}),$
where $u_{j}=F_{Y_{j}|\mathbf{X}}(y_{j}|\mathbf{x})$ are uniform marginal CDFs. The copula $C$ is a multivariate CDF on $[0,1]^{J}$ with uniform marginals. Our hybrid approach directly relates to this. The first step ($H_{1}$) models marginals $F_{Y_{j}|\mathbf{X}}$ and transforms them to the base distribution $F_{Z}$. Applying the probability integral transform (PIT), $u_{j}=F_{Z}(z_{1j})$, yields uniform marginals.
The copula density $c$ of $\bm{u}=(u_{1},\dots,u_{J})^{\top}$ is the ratio of the joint density and the product of the marginal densities:
$c(\bm{u}|\mathbf{x})=\frac{f_{\mathbf{Y}|\mathbf{X}}(F^{-1}_{Y_{1}|\mathbf{X}}(u_{1}|\mathbf{x}),\dots,F^{-1}_{Y_{J}|\mathbf{X}}(u_{J}|\mathbf{x})|\mathbf{x})}{\prod_{j=1}^{J}f_{Y_{j}|\mathbf{X}}(F^{-1}_{Y_{j}|\mathbf{X}}(u_{j}|\mathbf{x})|\mathbf{x})}.$
4 Numerical Experiments
Next, we evaluate our method using simulated and real-world datasets. All models are implemented using TensorFlow (v2.15.1) and TensorFlow Probability (v0.23.0) and trained using Adam (Tensorflow; TensorflowProbability; Kingma2017a). In our experiments, hyperparameters were chosen based on our previous experience and settings suggested in related works. However, we did some automated hyperparameter tuning (using Optuna) on some models applied to the simulated 2D data with performance monitored on a held-out validation set. This was mainly done to understand the influence of certain parameters, but it was not necessary to obtain a good model for each data set. We found that high-order Bernstein polynomials are needed for some datasets, suggesting the fixed knot placement of Bernstein polynomials could limit their flexibility(see Hothorn2018, for an ablation to the sensitivity of the Bernstein order).
Splines, on the other hand, might be more suitable for representing highly nonlinear or discontinuous functions, leading to a better fit on certain datasets. All hyperparameters used to generate the results presented are documented in the supplementary material and in the GitHub repository [^3] https://github.com/MArpogaus/hybrid_flows. This includes the code to perform the hyperparameter tuning along with the defined search spaces.
4.1 Benchmark Datasets
We evaluate our method on five common benchmark datasets POWER, GAS, HEPMASS, MINIBOONE, and BSDS300 (see Appendix Section B.5). We follow the preprocessing steps as in Papamakarios2018 and compare a MAF and our proposed hybrid MAF (HMAF). Both models utilize masked autoregressive transformations. Each MAF layer uses a masked affine transformation, parameterized by a masked neural network followed by an invertible linear transformation (1x1 convolution) initialized with a random (non-trainable) permutation. For the HMAF, we use a marginal transformation step with Bernstein polynomials before the MAFs layers.
For both MAFs and HMAF, we employ Rational Quadratic Splines (RQS) as the transformation functions within the MAFs layers of $H_{2}$. Hyperparameters, including the number of MAFs layers, the number of hidden units in the MADE networks, and the number of bins for the RQS transformations, were chosen based on the values reported in Durkan2019 and are detailed in the supplementary material and the GitHub repository. All models were trained using the Adam optimizer with early stopping after $50$ epochs without improvements and a cosine learning rate schedule.
Table 1: Test negative log-likelihood comparison against the state-of-the-art on real-world datasets (lower is better).
Values are averaged over 20 trials, and their spread is reported as two standard deviations.
model dataset name bsds300 gas hepmass miniboone power
HMAF test loss -153.663 $\pm$ 0.037 -11.625 $\pm$ 0.209 18.103 $\pm$ 0.058 12.057 $\pm$ 0.092 -0.527 $\pm$ 0.003
train loss -165.274 $\pm$ 0.124 -11.790 $\pm$ 0.230 17.856 $\pm$ 0.063 8.715 $\pm$ 0.415 -0.573 $\pm$ 0.004
validation loss -168.769 $\pm$ 0.035 -11.621 $\pm$ 0.210 18.123 $\pm$ 0.057 11.519 $\pm$ 0.070 -0.537 $\pm$ 0.003
MAF test loss -155.057 $\pm$ 0.065 -11.781 $\pm$ 0.032 18.090 $\pm$ 0.039 12.030 $\pm$ 0.073 -0.541 $\pm$ 0.003
train loss -167.694 $\pm$ 0.293 -11.881 $\pm$ 0.032 17.868 $\pm$ 0.041 9.605 $\pm$ 0.048 -0.606 $\pm$ 0.003
validation loss -170.188 $\pm$ 0.085 -11.779 $\pm$ 0.031 18.102 $\pm$ 0.040 11.546 $\pm$ 0.048 -0.551 $\pm$ 0.003
Table 1 presents the test log-likelihoods achieved by our models. Overall, HMAF demonstrates competitive performance compared to MAFs. Further investigation into more flexible marginal transformations could lead to even better performance and broader applicability of hybrid flow-based models. See Section B.5 for a visualization of individual model runs.
4.2 Simulated Data
We start with the classical bivariate (i.e. $J=2$) NFs datasets, moons and circles, exhibiting non-linear dependencies. Each dataset has $16,384$ data points (we reserved 25% for validation) generated using scikit-learn (scikit-learn). A binary feature $x$ based on spatial location is introduced. For moons, $x=1$ indicates the lower-right moon, and $x=0$ the upper-left. For circles, $x=1$ denotes the inner circle, and $x=0$ the outer. This assesses the models’ ability to capture $f(\mathbf{y}|x)$. We compare the following models:
  • multivariate normal (MVN) Assumes a conditional multivariate normal distribution parameterized by $x$ using a connected network (two layers, 16 units, ReLU activation).
  • multivariate conditional transformation model (MCTM): Employs Bernstein polynomials of order $M=300$ for marginal transformations and a triangular matrix $\Lambda$ for the dependency structure. In the conditional case, the marginal shift parameter $\beta_{x}$ and $\Lambda_{x}$ are obtained using Bernstein polynomials of order 6, as in Klein2022. While this approach offers interpretability through SAP, the linear dependence structure constrains its capacity.
  • coupling flow (CF)and masked autoregressive flow (MAF): These consist of two stacked coupling or masked autoregressive layers, parameterized by fully connected or masked neural networks, respectively (tree layers, 128 units, ReLU activation). Conditional models incorporate $x$ as input to an FCN whose output is added to the first layer’s parameters. We use Bernstein polynomials of order $M=300$ (CF (B), MAF (B)) and quadratic splines with 32 bins (CF (S), MAF (S)) as transformation functions.
  • hybrid CF (HCF): Combines elementwise Bernstein polynomials of order $M=300$ for marginals (similar to MCTMs) with a single coupling layer to model dependencies. Again Bernstein polynomials (B) and quadratic splines (S) are compared as transformation functions in the coupling layer. As the data exhibits a very complex distribution, simple shift terms are not enough to model the feature effect in the marginals. Instead, class-specific coefficients are used to model the feature effect on the Bernstein coefficients.
All models use the Adam optimizer (Kingma2017a) with early stopping and cosine learning rate decay (Loshchilov2017). All chosen hyperparameters are reported in the supplementary material.
Uncond. Cond. Uncond. Cond.
MVN
Refer to caption Refer to caption Refer to caption Refer to caption
MCTM
Refer to caption Refer to caption Refer to caption Refer to caption
CF (S)
Refer to caption Refer to caption Refer to caption Refer to caption
CF (B)
Refer to caption Refer to caption Refer to caption Refer to caption
MAF (S)
Refer to caption Refer to caption Refer to caption Refer to caption
MAF (B)
Refer to caption Refer to caption Refer to caption Refer to caption
HCF (S)
Refer to caption Refer to caption Refer to caption Refer to caption
HCF (B)
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 5: Estimated densities from different models fitted to simulated 2D data. Left two columns: Circles dataset, Right two columns: Moons dataset. Each cell shows the estimated density function, conditioned on $x$ (left) and not conditioned on $x$ (right). The conditional variable $x$ is 1 for the inner circle and the lower right moon, respectively.
Results
Figure 5 visualizes the estimated densities for different models on both the circles and moons datasets. Each cell displays contour plots of the estimated densities, both conditional (left) and unconditional (right). As expected, the multivariate normal (MVN) and the MCTM struggle with the non-linear dependencies. The MVN is restricted to elliptical distributions, while the MCTM, despite capturing marginals well, fails to represent the complex joint distribution due to its linear dependence structure. In contrast, CFs, MAFs, and hybrid CF (HCF) successfully capture the datasets’ non-linear shapes, demonstrating their ability to model complex distributions. Interestingly, models using Bernstein polynomials (CF (B), MAF (B), HCF (B)) tend to produce slightly blurrier results, indicating a bias towards smoother distributions compared to the spline-based models. Among the flexible models, quadratic spline transformations generally show a slight performance advantage over Bernstein polynomials for the CFs and MAFs architectures in terms of capturing fine-grained details of the data distribution, although the HCF models seem to perform equally well with both transformations.
Table 3 (in Section B.1) provides a quantitative comparison, reporting the average test negative logarithmic likelihood (NLL) across 20 trails. The results corroborate the visual observations from Figure 5. MVN and MCTM have significantly higher NLL values, indicating their inadequacy for these complex distributions. CFs and MAFs, especially with spline transformations, consistently achieve lower NLL. This difference underscores the importance of flexible dependency modeling. The HCF models perform well and demonstrate the effectiveness of the hybrid approach. The differences of Bernstein polynomial models in Figure 5 are directly reflected in lower NLL values, suggesting that these models lack certain flexibility to represent the data distribution. Consistent with the visualizations, including feature information improves model performance across all models, leading to consistently lower NLL values in the conditional setting. This improvement highlights the models’ ability to leverage the feature to better capture the conditional structure of the data.
4.3 Malnutrition Data
Finally, we evaluate our approach using a real-world dataset on childhood malnutrition in India, with three anthropometric indices (stunting, wasting, and underweight) as response variables. The goal is to model the joint distribution conditional on the child’s age (cage). We follow the data preprocessing steps from Klein2022.
We fit three models, all using Bernstein polynomials of order $M=6$ for the marginal transformations with a linear feature shift.
  • MCTM: Identical to the model specification in Klein2022, this model combines marginal transformations with a triangular matrix $\Lambda$. We capture the influence of cage on both the marginals (using linear shifts) and on the elements of $\Lambda$ using Bernstein polynomials of order $M=6$, i.e. the marginal shift is modeled as Bernstein polynomial $\bm{\beta}_{x}=\alpha(\text{cage})^{\top}\bm{\vartheta}$, where $\alpha$ represents the Bernstein basis, and $\bm{\vartheta}$ are the coefficients.
  • hybrid MAF (HMAF): For our approach, the MAFs for $H_{2}$ uses either Bernstein polynomials (B) or quadratic splines (S). We apply $H_{2}$ to $\mathbf{y}_{j>1}$, as $y_{1}$ is already normalized by $H_{1}$. Each MAFs layer is parameterized by a masked neural network conditional on cage. To capture dependencies between $y_{1}$ and $\mathbf{y}_{j>1}$, an additional fully connected neural network is used to condition the output of the MADE networks on $y_{1}$.
Results
Table 2 presents average test NLL. HMAF variants, especially HMAF (S), outperform MCTM, indicating that the non-linear dependence modeling of the MAFs is crucial. Additionally, this is confirmed by Figure 10 in Section B.2 via scatter, pairwise density, and marginal density plots of the dataset as well as samples drawn from the three models. The plots indicate that the MCTM captures marginals well, but fails to model dependencies. The HMAF model shows a greater resemblance to the observed data in the pairwise density and scatter plots.
Table 2: Average test NLL on the malnutrition data. Lower values indicate better performance. Values are averaged over 20 trials, and their spread is reported as two standard deviations.
model test loss
MCTM 3.470 $\pm$ 0.026
HMAF (S) 0.687 $\pm$ 0.206
HMAF (B) 2.062 $\pm$ 0.038
Marginal distributions
To assess whether the model captures the marginal distribution, Hothorn2014 suggest using quantile-quantile (Q-Q) plots to verify $\mathbf{W}$ follows the base distribution $F_{W}=\mathcal{N}(0,1)$. Figure 6 shows the empirical quantile function of marginally normalized samples $\mathbf{W}$ from the validation set plotted against the quantile function of a standard normal distribution evaluated at 200 equidistant probabilities. The results suggest a good fit of the transformation $H_{1}$, with slight underestimations of the marginals.
Figure 6: Q-Q plots of transformed samples against a standard normal distribution. Deviations from the diagonal indicate non-normality. The solid line represents the mean, while the shaded area indicates the 95% probability intervals obtained from 20 trials of randomly initialized models.
To assess the influence of $H_{2}$ on the marginals, we plotted the empirical quantiles of the dataset against an equal number of samples from all three models in Figure 7. All three models leave the normalized samples for stunting unchanged while introducing more deviations for wasting and underweight. This deviation is especially large for the MCTM models. The HMAF, particularly the spline variants, generally perform well but struggle with upper tails.
Figure 7: Q-Q plot comparing empirical quantiles of the dataset with those generated by the three models. The solid lines represent the mean, while the shaded areas indicate the 95% probability intervals obtained from 20 trials of randomly initialized models.
4.4 Interpretability
A crucial aspect of our proposed hybrid approach is its ability to model marginal distributions in an interpretable manner while simultaneously capturing complex dependencies. Using the Malnutrition data, we demonstrate the model’s interpretability in understanding learned relationships.
Feature-induced Changes in Marginal Distribution
Figure 8 shows the estimated marginal CDFs $F(y_{j}|\text{cage})$ and PDFs $f(y_{j}|\text{cage})$. The plots indicate non-linear shifts towards lower values as the age of the child increases indicating a deteriorating nutrition status according to all three malnutrition indicators. This overall trend can be explained by the fact that most children are born with a close-to-normal nutritional status but the effects of resource-scarce environments become increasingly relevant as children age.
Figure 8: Comparison of marginal effect on the PDFs and CDFs of stunting, wasting, and underweight with respect to cage.
Figure 9: Inverse marginal shift $-\bm{\beta}_{x}$. The plot demonstrates the relationship between marginal shifts and cage. The solid line represents the mean and the shaded area the 95% probability intervals over 20 trials of randomly initialized models.
Inverse Marginal Shift Terms
A more nuanced interpretation can be achieved by depicting the inverse marginal shift terms $\bm{\beta}_{x}$ as in Figure 9.
The plot shows the complex non-linear change in the nutritional status with increasing age, but also highlights that the change is more pronounced for stunting and underweight. This coincides with the intuition that acute malnutrition (as measured by the stunting indicator) materializes more quickly than chronic malnutrition (as measured by the wasting indicator). Underweight represents a mixture of both acute and chronic malnutrition, which again fits with the estimated shift term.
5 Discussion
We introduce a hybrid approach for density regression that combines the strengths of MCTMs and autoregressive NFs. In the first step of our HMAF approach, we model the marginal distributions using a transformation model with interpretable structured additive predictor. This enables transparent modeling of feature effects on the marginal distribution. In the second step, we use an autoregressive NF to effectively capture complex, non-linear dependencies between the response variables while preserving the marginals. Our results on simulated and real-world datasets showed that our method is competitive with state-of-the-art methods.
The hybrid approach offers 1) interpretability, by transparently modeling of feature effects using structured additive predictors, 2) flexibility, by capturing complex non-linear dependencies using an autoregressive flow, and 3) efficiency, by offering fast computation of log-likelihoods and gradients due to the MADE architecture.
Acknowledgements.
During the finalization of this paper, large language models (LLMs) were used to optimize language, and grammar, and restructure some parts of the document. This research was funded by the Carl-Zeiss-Stiftung in the project ”DeepCarbPlanner” (grant no. P2021-08-007). We thank the DACHS data analysis cluster, hosted at Hochschule Esslingen and co-funded by the MWK within the DFG’s ,,Großgeräte der Länder” program, for providing the computational resources necessary for this research [^4] https://wiki.bwhpc.de/e/DACHS.
[…]
Hybrid Bernstein Normalizing Flows for Flexible Multivariate Density Regression with Interpretable Marginal
(Supplementary Material)
Appendix A Extended Background
A.1 Bernstein Polynomials
Bernstein polynomials of order $M$ are defined as
$h(z)=\frac{1}{M+1}\sum_{i=0}^{M}\operatorname{Be}_{i}(z)\vartheta_{i},$ (2)
where $\operatorname{Be}_{i}(z)$ is the density of a Beta distribution with parameters $i+1$ and $M-i+1$, and $\vartheta_{0},\ldots,\vartheta_{M}$ are the Bernstein coefficients Farouki2012. As $M$ increases, Bernstein polynomials become a universal approximator for smooth functions in $[0,1]$ Farouki1988. In practice, $M\geq 10$ is often sufficient Hothorn2018.
Bernstein polynomials are defined for values of $z$ within the range $[0,1]$. Outside this interval, linear extrapolation is performed. To guarantee invertibility, transformation functions must be bijective, achieved by strict monotonicity. When using Bernstein polynomials, monotonicity is enforced by constraining the Bernstein coefficients $\vartheta_{0},\ldots,\vartheta_{M}$ to be increasing. This is achieved by recursively applying a strictly positive function like $\operatorname{softplus}$ to an unconstrained vector $\tilde{\vartheta}_{0},\ldots,\tilde{\vartheta}_{M}$, such that $\vartheta_{0}=\tilde{\vartheta}_{0}$ and $\vartheta_{k}=\vartheta_{k-1}+\operatorname{softplus}(\tilde{\vartheta}_{k})$ for $k=1,\ldots,M$. Parameter estimation can depend on initialization, as $\vartheta_{0}$ is directly derived from unconstrained parameters. Monotonicity of $h$ is enforced by constraining the coefficients to be increasing, e.g. by $\vartheta_{0}=\tilde{\vartheta}_{0}$ and $\vartheta_{k}=\vartheta_{k-1}+\operatorname{softplus}(\tilde{\vartheta}_{k})$ for $k=1,\ldots,M$ Sick2021. The inverse transformation can be found using a root-finding algorithm Chandrupatla1997.
We require transformations to cover at least the range $[-3,3]$ (e.g., ±$\pm 3\sigma$ of a standard Gaussian). Since Bernstein polynomials’ boundaries are defined by its first and last coefficients ($f(0)=\vartheta_{0}$ and $f(1)=\vartheta_{M}$), we determine these from unrestricted parameters $\tilde{\vartheta}_{0}$ and $\tilde{\vartheta}_{M+1}$ via $\vartheta_{0}=-\operatorname{softplus}(\tilde{\vartheta}_{0})-3.0\leq-3$ and $\vartheta_{M}=\operatorname{softplus}(\tilde{\vartheta}_{M+1})+3.0\geq 3$. To ensure $\sum_{k=1}^{M}{(\vartheta_{k}-\vartheta_{k-1})}=\vartheta_{M}-\vartheta_{0}=:\Delta$, the remaining coefficients $\vartheta_{k}$ for $k=1,\ldots,M$ are calculated as:
$\vartheta_{k}=\vartheta_{k-1}+\Delta\cdot\operatorname{softmax}\left(\left[\tilde{\vartheta}_{1},\tilde{\vartheta}_{3},\ldots,\tilde{\vartheta}_{M}\right]\right)_{k-1}$ (3)
Since $\Delta$ and all $\operatorname{softmax}$ components are non-negative, $\vartheta_{k}-\vartheta_{k-1}\geq 0$, ensuring monotonicity.
A.2 Maximum Likelihood Estimation for Parameter Inference
Parameter estimation usually involves minimizing a divergence between the true distribution $p_{y}(y)$ and the transformation model $p_{y}(\mathbf{y}|\theta)$ regarding the model’s parameters $\theta$ (Papamakarios2021). Since $p_{y}(y)$ is unknown, we minimize the negative log-likelihood of the empirical distribution $p_{\mathcal{D}}$ obtained from a finite set $\mathcal{D}$ of $N$ i.i.d. observations $\mathbb{Y}=\{\mathbf{y}_{1},\ldots,\mathbf{y}_{m}\}$:
$\operatorname*{NLL}=-\sum_{y\in\mathcal{D}}\log\left(p_{y}(y;\theta)\right)$ (4)
which is equivalent to minimizing the KL divergence between $p_{\mathcal{D}}$ and the flow-based model $p_{y}(y;\theta)$ (Papamakarios2021).
A.3 Multivariate Conditional Transformation Models
MCTM (Klein2022) use element-wise transformations $\tilde{h}_{j}\left(y_{j}\right),j=1,\ldots,J$ and a linear triangular $(J\times J)$ matrix $\Lambda$ to model correlations:
$h_{j}(y_{1},\ldots,y_{j})=\lambda_{j1}\tilde{h}_{1}(y_{1})+\ldots+\lambda_{j,j-1}\tilde{h}_{j-1}(y_{j-1})+\tilde{h}_{j}(y_{j})$ (5)
For conditional distributions, $\Lambda$ and transformation parameters $\theta$ can depend on covariates $\mathbf{x}$:
$h_{j}(\mathbf{y}|\mathbf{x})=\sum_{i=1}^{j-1}\lambda_{ji}(\mathbf{x})\tilde{h}_{i}(y_{i};\theta_{i}(\mathbf{x}))+\tilde{h}_{j}(y_{j};\theta_{j}(\mathbf{x}))$ (6)
A.4 Autoregressive Transformation Models
Autoregressive flows factorize multivariate distributions based on the chain rule of probability:
$p_{y}(\mathbf{y})=\prod_{i=1}^{D}p_{y}(y_{i}|\mathbf{y}_{<i})$ (7)
Applying the change of variables formula yields:
$p_{y}(\mathbf{y})=p_{z}\left(h(y)\right)\left|\det\nabla{h}(y)\right|=\prod_{i=1}^{D}p_{z}\left(h_{i}(y_{i},\mathbf{y}_{<i})\right)\left|\det\nabla{h}_{i}(y_{i},\mathbf{y}_{<i})\right|$ (8)
where $h_{i}$ is a diffeomorphism applied to the $i$-th element of $\mathbf{y}$, conditioned on preceding elements $\mathbf{y}_{<i}$. The lower triangular Jacobian determinant is the product of its diagonal elements:
$\det\nabla{h}(y)=\prod_{i=1}^{D}\frac{\partial F_{i}}{\partial y_{i}}$ (9)
This is implemented as $z_{i}=h_{i}(y_{i},\theta_{i})$ with $\theta_{i}=c(\mathbf{y}_{<i})$, where the conditioner $c_{i}$, often a neural network, captures dependencies and enforces the autoregressive property (Papamakarios2021). New samples are obtained via the inverse transformation $y_{i}=T_{i}^{-1}(z_{i},\theta_{i})$.
Two common architectures for the conditioner are coupling layers and masked autoregressive networks. Coupling flows (Dinh2017) split the response $\mathbf{y}\in\mathbb{R}^{D}$ into subsets $(\mathbf{y}_{A},\mathbf{y}_{B})\in(\mathbb{R}^{d},\mathbb{R}^{D-d})$, with one subset conditioning the transformation of the other:
$\mathbf{y}=\begin{cases}\mathbf{y}_{A}=\mathbf{y}_{A}\\ \mathbf{y}_{B}=h(\mathbf{y}_{B};\Theta(\mathbf{y}_{A}))\end{cases}$ (10)
MADE (Germain2015), as used in MAFs (Papamakarios2018) and inverse autoregressive flows (IAFs) (Kingma2016), generalizes this using masked neural networks to enforce autoregressive constraints.
In MADE, the output $d$ depends only on inputs $\mathbf{y}_{<d}$. This is achieved by element-wise multiplying weight matrices by binary masks, zeroing connections that violate the autoregressive property. Whether the masked network uses $\mathbf{y}$ or the latent representation $\mathbf{z}$ as input affects only whether inference or sampling is iterative (Papamakarios2021).
A.5 Structured Additive Predictors for Enhanced Flexibility
Structured Additive Predictors (SAP; Fahrmeir2013; Fahrmeir2009) allow for (non-)linear effects, interactions, and other structured terms. An exemplary predictor is given by
$\eta(\mathbf{x})=\beta_{0}+\sum_{u=1}^{U}f_{u}(x_{u})+\sum_{u<v}f_{uv}(x_{u},x_{v})+\dots$ (11)
Here, $f_{u}(x_{u})$ is a linear or non-linear function, where the latter is usually specified using regression splines to stay in the context of parametric regression. $f_{uv}(x_{u},x_{v})$ are linear or smooth interaction effects. The order of interaction determines the degree of interpretability. In our hybrid approach, SAP can be used to parameterize marginal shifts $\beta_{j}(\mathbf{x})$ in $H_{1}$.
Appendix B Extended Results
B.1 negative log-likelihood on 2D simluation datasets
dataset name circles moons
conditional False True False True
model
MVN -0.204 $\pm$ $0$0.423 $\pm$ 0.002 -0.151 $\pm$ 0.002 -0.704 $\pm$ 0.004
MCTM -0.490 $\pm$ 0.002 -0.489 $\pm$ 0.002 -0.536 $\pm$ $0$1.046 $\pm$ 0.006
MAF (S) -1.123 $\pm$ 0.022 -1.123 $\pm$ 0.022 -1.611 $\pm$ 0.042 -1.611 $\pm$ 0.042
MAF (B) -1.179 $\pm$ 0.014 -1.179 $\pm$ 0.014 -1.625 $\pm$ 0.016 -1.625 $\pm$ 0.016
CF (S) -1.045 $\pm$ 0.132 -1.861 $\pm$ 0.056 -1.587 $\pm$ 0.052 -2.306 $\pm$ 0.050
CF (B) -0.651 $\pm$ 0.202 -1.657 $\pm$ 0.038 -1.350 $\pm$ 0.106 -2.186 $\pm$ 0.116
HCF (S) -1.175 $\pm$ 0.012 -1.870 $\pm$ 0.018 -1.628 $\pm$ 0.018 -2.332 $\pm$ 0.014
HCF (B) -1.071 $\pm$ 0.024 -1.826 $\pm$ 0.078 -1.583 $\pm$ 0.042 -2.332 $\pm$ 0.032
Table 3: Test negative log-likelihood on 2D simluation datasets (lower is better). Log-likelihoods are averaged over 20 trials, and their spread is reported as two standard deviations.
B.2 Scatter plots of Samples from Malnutrition Models
Figure 10 shows that MCTM captures marginals well but fails to model dependencies. The HMAF models, especially with spline transformations (HMAF (S)), show a greater resemblance to the observed data in the pairwise density plots.
(a) Random samples from the validation dataset.
(b) Random samples from the MCTM model.
(c) Random samples from the hybrid model using quadratic splines as the transformation function.
(d) Random samples from the hybrid model using Bernstein polynomials as the transformation function.
Figure 10: Scatter plots of the three target variables—stunting, wasting, and underweight. In the upper triangular section, random data points are illustrated, either from the validation dataset or sampled from the three models. The diagonal displays the marginal kernel density estimation (KDE) plots, while the lower triangular region contains the two-dimensional KDE plots. The data is categorized by the variable cage.
B.3 Interpretability of Covariate Effects
Interpretable covariate effects are important in many regression tasks. CTMs offer a balance between interpretability and flexibility (Hothorn2014). The base distribution and transformation function parameterization determine the interpretability.
In a univariate CTM with covariate $x$:
$\mathbb{P}(Y\leq y|X=x)=F_{Z}(h(y|\theta_{x})),$ (12)
where $h(y|\theta_{x})=\alpha(y)^{\top}\bm{\vartheta}+\beta x$, with $\beta$ is the covariate effect on the transformed response, and $\alpha(y)^{\top}\bm{\vartheta}$ controlling the distribution’s shape.
  • Logistic Distribution: $\beta$ is the log-odds ratio, quantifying the change in the odds of $Y\leq y$ associated with a unit increase in $x$.
  • Minimum Extreme Value Distribution: $\beta$ is the log-hazard ratio, representing the influence of $x$ on the instantaneous risk of an event.
  • Gaussian Distribution: With a linear transformation $h(y|\theta_{x})$, $\beta$ is the change in $Y$’s conditional mean per unit change in $x$ (scaled by the standard deviation). With non-linear $h$, $\beta$’s interpretation is less direct, affecting multiple moments.
B.4 Interpretable Covariate Effect on the Dependency Structure in MCTMs
As described in Klein2022, the covariate effect on the dependency in the MCTM model can be interpreted as Spearman’s rank correlation from the covariance matrix
$\Sigma=\Lambda^{\top}\Lambda^{-\top}$ (13)
via the Pearson correlation coefficient,
$\rho_{ij}=\frac{\Sigma_{ij}}{\sigma_{i}\sigma_{j}}=\frac{\Sigma_{ij}}{\sqrt{\Sigma_{ii}\Sigma_{jj}}}$ (14)
where $\Sigma_{ij}$ represents the covariance between variables $y_{i}$ and $y_{j}$ and $\Sigma_{ii}$ the variance $\sigma_{i}^{2}$. To convert Pearson correlations to Spearman’s rank correlation, the following transformation is applied:
$\rho_{s}=\frac{6}{\pi}\arcsin\left(\frac{\rho}{2}\right)$ (15)
The resulting covariate-dependent rank correlations are shown in Figure 11.
Figure 11: Comparison of Spearman’s rank correlation $\rho^{S}$ estimates between stunting, wasting, and underweight with respect to cage.
B.5 Benchmark Datasets
We evaluate our method on five common benchmark datasets: POWER [^5] https://archive.ics.uci.edu/dataset/235, GAS [^6] https://archive.ics.uci.edu/dataset/322. HEPMASS [^7] https://archive.ics.uci.edu/dataset/347, MINIBOONE [^8] https://archive.ics.uci.edu/dataset/199/, BSDS300999https://www2.eecs.berkeley.edu/Research/Projects/CS/vision/bsds/. We follow the preprocessing steps from Papamakarios2018101010https://github.com/francois-rozet/uci-datasets.
Figure 12 shows the distribution of NLL scores for HMAF and MAF across 20 different random initializations. The figure visualizes performance robustness. Consistent lower NLL values indicate better performance, smaller IQRs suggest less sensitivity to initialization, and a lower spread of outliers suggests more stable training.
Figure 12: Distribution of negative log-likelihood (NLL) scores for the HMAF and MAF models, resulting from 20 runs with different random weight initializations. The box plots show the median, interquartile range (IQR), and the data range of 1.5 times the IQR, outside the IQR. The swarm plots above the boxes show the individual NLL scores for each run.
Appendix C Complexity Estimates (Runtime and Number of Parameters)
C.1 Runtime of Model Variants
The tables below present the runtime for training and evaluating our models on the HPC Cluster at the University of Applied Sciences Esslingen, utilizing NVIDIA L40S GPUs with 48 GB VRAM. Note that our code is not fully optimized for runtime, and variations in training and inference times are significantly influenced by hyperparameters such as the number of epochs and the use of early stopping.
Table 4: Runtime in Minutes for training and evaluation of models on benchmark data. Variance resulting deviations from 20 runs reported as standard deviation.
model dataset name train evaluation
HMAF bsds300 1191.992 $\pm$ 537.944 481.991 $\pm$ 0.650
gas 319.809 $\pm$ 132.472 14.931 $\pm$ 0.043
hepmass 229.736 $\pm$ 155.484 15.094 $\pm$ 0.047
miniboone 82.882 $\pm$ 58.692 3.933 $\pm$ 0.012
power 437.108 $\pm$ 63.707 11.321 $\pm$ 0.091
MAF bsds300 261.977 $\pm$ 70.957 16.716 $\pm$ 0.022
gas 68.993 $\pm$ 0.073 1.858 $\pm$ 0.005
hepmass 34.774 $\pm$ 0.003 1.540 $\pm$ 0.004
miniboone 16.486 $\pm$ 1.404 0.279 $\pm$ 0.001
power 136.796 $\pm$ 0.004 4.979 $\pm$ 0.120
Table 5: Mean runtime in seconds for training and evaluation of models on simulated data. Variance resulting deviations from 20 runs reported as standard deviation.
train evaluation
dataset name conditional model
circles False CF (B) 35.043 $\pm$ 15.008 16.782 $\pm$ 0.238
CF (S) 317.228 $\pm$ 233.129 17.428 $\pm$ 0.181
HCF (B) 110.039 $\pm$ 11.230 57.759 $\pm$ 0.830
HCF (S) 76.960 $\pm$ 37.909 29.489 $\pm$ 0.191
MAF (B) 91.213 $\pm$ 0.354 19.562 $\pm$ 2.876
MAF (S) 145.815 $\pm$ 36.820 17.369 $\pm$ 1.585
MCTM 42.597 $\pm$ 0.724 29.009 $\pm$ 0.375
MVN 44.892 $\pm$ 0.099 10.722 $\pm$ 0.067
True CF (B) 41.655 $\pm$ 20.906 20.423 $\pm$ 0.223
CF (S) 329.650 $\pm$ 211.377 17.864 $\pm$ 0.267
HCF (B) 66.384 $\pm$ 40.931 73.782 $\pm$ 1.620
HCF (S) 171.494 $\pm$ 107.205 31.913 $\pm$ 0.128
MAF (B) 91.153 $\pm$ 0.337 20.306 $\pm$ 2.840
MAF (S) 145.765 $\pm$ 36.843 17.912 $\pm$ 1.554
MCTM 304.875 $\pm$ 112.109 31.375 $\pm$ 2.577
MVN 56.360 $\pm$ 36.401 11.708 $\pm$ 0.055
moons False CF (B) 33.991 $\pm$ 12.603 16.842 $\pm$ 0.145
CF (S) 443.163 $\pm$ 80.343 17.129 $\pm$ 0.312
HCF (B) 184.718 $\pm$ 71.683 67.237 $\pm$ 2.418
HCF (S) 76.214 $\pm$ 39.760 29.368 $\pm$ 0.106
MAF (B) 91.252 $\pm$ 0.538 19.872 $\pm$ 3.607
MAF (S) 134.682 $\pm$ 62.895 14.719 $\pm$ 3.458
MCTM 42.544 $\pm$ 0.742 28.949 $\pm$ 0.141
MVN 44.823 $\pm$ 0.024 11.497 $\pm$ 0.046
True CF (B) 44.995 $\pm$ 20.047 20.477 $\pm$ 0.313
CF (S) 368.668 $\pm$ 250.469 17.644 $\pm$ 0.415
HCF (B) 48.243 $\pm$ 17.635 40.735 $\pm$ 0.516
HCF (S) 169.016 $\pm$ 62.618 31.744 $\pm$ 0.146
MAF (B) 91.215 $\pm$ 0.529 20.623 $\pm$ 3.563
MAF (S) 135.603 $\pm$ 68.688 15.308 $\pm$ 3.487
MCTM 475.688 $\pm$ 1.217 35.501 $\pm$ 7.524
MVN 48.118 $\pm$ 18.322 11.030 $\pm$ 0.071
Table 6: Mean runtime in seconds for training and evaluation of models on malnutrition data. Variance resulting deviations from 20 runs reported as standard deviation.
train evaluation
model
HMAF (B) 260.752 $\pm$ 121.895 20.823 $\pm$ 0.535
HMAF (S) 1993.317 $\pm$ 717.933 19.649 $\pm$ 0.110
MCTM 4106.187 $\pm$ 725.136 16.877 $\pm$ 0.847